OpenStudy (anonymous):

evaluate the expression if X = 3 and Y = 11 3x + 2y

Directrix (directrix):

3x + 2y X = 3 and Y = 11 3*3 + 2*11 = ? The idea is to substitute 3 for the variable x and 11 for the variable y and then evaluate the resulting expression.

Directrix (directrix):

3*3 + 2*11 = ? Post what you get, okay?

OpenStudy (anonymous):

31

Directrix (directrix):

Correct: 9 + 22